OpenStudy (anonymous):

5x > 35 Answers to choose from x = 7 x > 7 x < 7 x ≥ -7 x > -7 I was thinking it would be x > 7, am I right?

OpenStudy (whpalmer4):

\[5x>35\]We can divide both sides of the inequality by 5, giving us: \[\frac{5x}5 > \frac{35}5\]or\[x>7\]Therefore, you are correct. You have to be careful when doing this as dividing (or multiplying) by a negative number reverses the direction of the inequality. For example, if we divided \(x>7\) by \(-1\), we get \[\frac{x}{-1} < \frac{7}{-1}\]\[-x<-7\]If you draw the number line and mark where each inequality is true, it shouldn't be hard to convince yourself that the two are equivalent.